OpenStudy (anonymous):

Which sequence is modeled by the graph below? coordinate plane showing the points 2, 8; 3, 4; 4, 2; and 5, 1

OpenStudy (anonymous):

will fan and medal!!! Answer choices are: an = one half(16)n − 1 an = 4(2)n − 1 an = 16(−2)n − 1 an = 16(one half)n − 1

OpenStudy (anonymous):

@bibby @Brainybeauty ???